Find the value of the variable that satisfies the equation. Check your solution. Answers that are not integers may be left in fractional or decimal form.

8x=48+2x

Answer by HyperBrain(694)

8x=48+2x
8x-2x=48+2x-2x
6x=48
x=48/6
x=8
Power up,
HyperBrain!
Stop this mess! Click here